OpenStudy (anonymous):

Math Help Please!!! An enclosed gas exerts a pressure P on the walls of a container. This pressure is directly proportional to the temperature T of the gas, measured in Kelvin. If the pressure is 7 lb per square inch when the temperature is 420K, find the constant of variation. A) 60 B) 1/60 C) 49 D) 1/49

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (godlovesme):

p=tk 7=420k solve for k

OpenStudy (anonymous):

It is either A or B because 420 divided by 7 is 60

OpenStudy (godlovesme):

it's 7/420 which reduces to 1/60 but lemme check though

OpenStudy (anonymous):

It's B.

OpenStudy (anonymous):

But how do you know? @ApoorvaAnand

OpenStudy (anonymous):

P = (k)T => 7 = (k)420 => k = 7/420 = 1/60 = B

OpenStudy (anonymous):

Thank you @ApoorvaAnand